OpenStudy (anonymous):

If a-3b=9b-7a then what is the ratio of a to b?

OpenStudy (anonymous):

start by using algebra to get all the a's on one side and all the b's on the other

OpenStudy (anonymous):

a-3b = 9b-7a 8a = 12b

OpenStudy (anonymous):

Then do the next step to transform 8a = 12b into an equation with the ratio (a/b) on the left... what will the number on the right side be?

OpenStudy (anonymous):

Would it be 8a/b = 12b

OpenStudy (anonymous):

not 12 b just 12 i mean

OpenStudy (anonymous):

you need to divide both sides by 8 to get rid of the 8 on the left side... 8(a/b) = 12 (a/b) = 12/8 then you should probably reduce that fraction...